OpenStudy (anonymous):

Basically i would like for someone to solve this double integral. sqrt(1-x2-y2)/y D: x=0 ; x2+y2=1 Tried to solve it, wolfram cant, please need help!

OpenStudy (lgbasallote):

could you draw it out?

OpenStudy (anonymous):

i used Trigonometric Substitution \[\int\limits_{?}^{?}\int\limits_{?}^{?}(\sqrt{1-x2-y2}/y)dxdy\] still doesn't work?